# Symmetric teleparallel gravitational effects on solar neutrino oscillations

### v1

@@ 1 Introduction
Neutrino oscillations, first proposed by Pontecorvo [1] and later confirmed by solar [2, 3], atmospheric [4], reactor [5, 6], and accelerator experiments [7, 8], stand among the most profound discoveries in modern particle physics. The mere fact that neutrinos change flavour during propagation unequivocally demonstrates their non-zero masses and that flavour eigenstates are quantum superpositions of mass eigenstates—a phenomenon squarely outside the original Standard Model.
@@
Einstein’s GR is a mathematically coherent theory that has achieved remarkable success in accounting for a broad spectrum of observational phenomena. Within this framework, gravity is geometrically interpreted through the curvature derived from the metric structure of Riemannian spacetime. Nevertheless, the difficulty of explaining dark matter and dark energy within the theory, together with the absence of a fully consistent quantization and related conceptual challenges, provides compelling motivation to investigate alternative gravitational models. Several avenues exist for modifying gravity, and generalized geometric frameworks extending beyond Riemannian spacetime present especially intriguing prospects.
@@ 2 Metric-affine spacetimes
where $\mathrm{d}$ is the exterior derivative and $D$ is the covariant exterior derivative. Here we used $\mathrm{d}\eta_{ab}=0$ in the first equation. The notation $(\bullet)$ and $\bullet$] denotes symmetrization and anti-symmetrization with respect to the enclosed indices: $\omega_{(ab)}=\frac{1}{2}(\omega_{ab}+\omega_{ba})$ and $\omega_{[ab]}=\frac{1}{2}(\omega_{ab}-\omega_{ba})$. The first kind trace (Weyl) $Q$ and independently the second kind trace $P$ 1-forms constructed from the non-metricity will be used in the forthcoming sections:
@@
Geometry is classified according to whether non-metricity, torsion, and/or curvature vanish. Accordingly, symmetric teleparallel geometry is defined by the vanishing of both curvature and torsion, $R^{a}{}_{b}=0$ and $T^{a}=0$, while non-metricity remains unrestricted, $Q_{ab}\neq 0$. The condition $R^{a}{}_{b}=0$ implies that the connection is flat, meaning that parallel transport is path-independent. Combined with $T^{a}=0$, this ensures the existence of global inertial frames where the connection can be made to vanish in the holonomic coordinate bases—the natural gauge or inertial gauge or coincident gauge [13],[18]. In the orthonormal bases, the coincident gauged connection 1-form takes the simple form
@@ 3 Dirac equation in metric-affine spacetimes
where the spinor connection 1-form is
@@
Here $\sigma_{ab}=\frac{1}{4}[\gamma_{a},\gamma_{b}]$ are the Lorentz generators, $\gamma_{a}$ are the Dirac matrices satisfying $\{\gamma_{a},\gamma_{b}\}=2\eta_{ab}I$, $\gamma_{5}=\gamma_{0}\gamma_{1}\gamma_{2}\gamma_{3}$, and $a_{1},a_{2},a_{3},a_{4},b_{3},b_{4}$} are complex coupling constants. The coupling constants must satisfy specific constraints for consistency between the variational and canonical Dirac equations [22]:
@@ 4 Neutrinos propagating at the equator of the Sun
Above, $t=0$ denotes the production time of the neutrino on the solar surface, while $t>0$ is the observation time on the Earth’s surface; thus $\nu^{\uparrow}(0)$ and $\nu^{\downarrow}(0)$ are integration constants representing the initial state of the neutrinos. Since neutrinos are ultra-relativistic particles, in the phase integrals one may use $p\approx E$ and $dt\approx dr$.
@@
In that case, what we need to do now is to compute the eigenvalues of the Dirac Hamiltonian. We shall accomplish this in two steps. First, we bring the matrix $H$ into block-diagonal form; then, we diagonalize the upper block. For the first step, we render the matrix $H$ block-diagonal by means of the unitary transformation matrix $U$ by changing $H\psi=E\psi$ to $H^{\prime}\psi^{\prime}=E\psi^{\prime}$ such that
$H^{\prime}=U^{\dagger}HU=\begin{pmatrix}H_{+}&0\\ 0&H_{-}\end{pmatrix}\quad\text{and}\quad\psi^{\prime}=U^{\dagger}\psi=\begin{pmatrix}\psi_{+}\\ \psi_{-}\end{pmatrix}$, (22)
where $H_{+}$ and $H_{-}$ are two-component square matrices, and $\psi_{+}$ and $\psi_{-}$ are two-component column matrices. Then, under certain assumptions, which we have made on the basis of observational data to be discussed later, $H_{+}$ has been approximately calculated as follows:
@@
$\mathcal{C}$ $:=a\left(\frac{f}{2r^{2}}-i\frac{f^{3}}{r}p+\frac{i\mathcal{M}f^{2}}{r}\right).$ (24c)
In the second step, we compute the eigenvalues of the Hamiltonian $H_{+}$ up to order $a$ by means of the eigenvalue equation $H_{+}\psi_{+}=E_{+}\psi_{+}$:
@@
$P_{\nu_{e}\to\nu_{\mu}}$ $=\sin^{2}(2\theta),\sin^{2}\left(\frac{\Delta\Phi}{2}\right)$ (29)
@@
  • Both mass eigenstates can be spin-up.
@@
  • Both mass eigenstates can be spin-down.
@@
  • $\nu_{2}$ can be spin-up while $\nu_{1}$ can be spin-down.
@@
  • $\nu_{2}$ can be spin-down while $\nu_{1}$ can be spin-up.
@@
    Spin-dependent geometric phase: The parameters $a_{2}$ and $a_{4}$ contribute only when the mass eigenstates have opposite helicities. This asymmetric contribution positive for one configuration, negative for the other averages to zero for unpolarized neutrino beams but could manifest in polarization-sensitive experiments.
@@ Acknowledgments
This study was supported by Scientific and Technological Research Council of Türkiye (TUBITAK) under Grant Number 124F325. The authors thank TUBITAK for their support.

### v2

@@ 1 Introduction
Neutrino oscillations, first proposed by Pontecorvo [1] and later confirmed by solar [2, 3], atmospheric [4], reactor [5, 6], and accelerator experiments [7, 8], stand among the most profound discoveries in modern particle physics. The mere fact that neutrinos change flavour during propagation unequivocally demonstrates their non-zero masses and that flavour eigenstates are quantum superpositions of mass eigenstates—a phenomenon squarely outside the original Standard Model. In solar neutrino oscillations, a particle that essentially leaves the solar surface as an electron neutrino is detected at the Earth’s surface as a muon neutrino. Therefore, the measured electron neutrino flux is smaller than expected.
@@
Einstein’s GR is a mathematically coherent theory that has achieved remarkable success in accounting for a broad spectrum of observational phenomena. Within this framework, gravity is geometrically interpreted through the curvature derived from the metric structure of Riemannian spacetime. Nevertheless, the difficulty of explaining dark matter and dark energy within the theory, together with the absence of a fully consistent quantization and related conceptual challenges, provides compelling motivation to investigate alternative gravitational models [15]. Several avenues exist for modifying gravity, and generalized geometric frameworks extending beyond Riemannian spacetime present especially intriguing prospects.
@@ 2 Metric-affine spacetimes
where $\mathrm{d}$ is the exterior derivative and $D$ is the covariant exterior derivative. Here we used $\mathrm{d}\eta_{ab}=0$ in the first equation. Throughout the paper, the notations $(\bullet)$ and $\bullet$] denote symmetrization and anti-symmetrization with respect to the enclosed indices: $\omega_{(ab)}=\frac{1}{2}(\omega_{ab}+\omega_{ba})$ and $\omega_{[ab]}=\frac{1}{2}(\omega_{ab}-\omega_{ba})$. The first kind trace (Weyl) $Q$ and independently the second kind trace $P$ 1-forms constructed from the non-metricity will be used in the forthcoming sections:
@@
Geometry is classified according to whether non-metricity, torsion, and/or curvature vanish. Accordingly, symmetric teleparallel geometry is defined by the vanishing of both curvature and torsion, $R^{a}{}_{b}=0$ and $T^{a}=0$, while non-metricity remains unrestricted, $Q_{ab}\neq 0$. The condition $R^{a}{}_{b}=0$ implies that the connection is flat, meaning that parallel transport is path-independent. Combined with $T^{a}=0$, this ensures the existence of global inertial frames where the connection can be made to vanish in the holonomic coordinate bases—the natural gauge or inertial gauge or coincident gauge [17, 22]. In the orthonormal bases, the coincident gauged connection 1-form takes the simple form
@@ 3 Dirac equation in metric-affine spacetimes
where $\psi$ is a Dirac spinor and the spinor connection 1-form is
@@
Here $\sigma_{ab}=\frac{1}{4}[\gamma_{a},\gamma_{b}]$ are the Lorentz generators, $\gamma_{a}$ are the Dirac matrices satisfying $\{\gamma_{a},\gamma_{b}\}=2\eta_{ab}I$, $\gamma_{5}=\gamma_{0}\gamma_{1}\gamma_{2}\gamma_{3}$, and $a_{1},a_{2},a_{3},a_{4},b_{3},b_{4}$} are complex coupling constants.
In Riemannian spacetime, only one of the basis elements of the Clifford algebra, namely $\sigma_{ab}$, appears in the spinor connection. However, if one requires that all basis elements of the Clifford algebra be present in the spinor connection in a metric-affine spacetime, this combination plus the term $(b_{1}I+b_{2}\gamma_{5})\,T$ emerge, where $T=\iota_{a}T^{a}$ is the torsion trace $1$-form, see [29]. We suspect that the new coupling constants encode traces of novel types of interactions of the spinor. For the time being, the physical interpretation of these constants remains an open problem. The present work has been carried out precisely with this aim. Meanwhile, the coupling constants must satisfy specific constraints for consistency between the variational and canonical Dirac equations:
@@ 4 Neutrinos propagating at the equator of the Sun
Above, $t=0$ denotes the production time of the neutrino on the solar surface, while $t>0$ is the observation time on the Earth’s surface; thus $\nu^{\uparrow}(0)$ and $\nu^{\downarrow}(0)$ are integration constants representing the initial state of the neutrinos. Since neutrinos are ultra-relativistic particles, in the phase integrals one may use $p\approx E$ and $dt\approx dr$:
@@
In that case, what we need to do now is to compute the eigenvalues of the Dirac Hamiltonian in Eq.(16) which is a complicated four-dimensional square matrix. We shall accomplish this in two steps. First, we bring the matrix $H$ into block-diagonal form; then, we diagonalize the upper block. For the first step, we perform the following calculation:
$H\psi=H_{\pm}\psi\qquad\Rightarrow\qquad\text{det}(H-H_{\pm}I)=\text{det}\begin{pmatrix}H_{11}-H_{\pm}&H_{12}\\ H_{21}&H_{22}-H_{\pm}\\ \end{pmatrix}=0\,$, (22a)
$H_{\pm}=\frac{1}{2}\left\{\left(H_{11}+H_{22}\right)\pm\sqrt{\left(H_{11}+H_{22}\right)^{2}-4\left(H_{11}H_{22}-H_{12}H_{21}\right)}\right\}\,$, (22b)
where $H_{11},\,H_{12},\,H_{21},\,H_{22}$ are the two dimensional submatrices of $H$. Now, the two-dimensional square matrices $H_{\pm}$ are the diagonal elements of $H$. Then, under certain assumptions based on the observational data to be discussed later, $H_{+}$ has been approximately calculated as follows:
@@
$\mathcal{C}$ $:=a\left(\frac{f}{2r^{2}}-i\frac{f^{3}}{r}p+\frac{i\mathcal{M}f^{2}}{r}\right),$ (24c)
where $\mathcal{M}:=m-4b_{3}$. In the second step, we compute the eigenvalues of the Hamiltonian $H_{+}$ up to order $a$ by means of the eigenvalue equation $H_{+}\psi_{+}=E_{+}\psi_{+}$:
@@
$P_{\nu_{e}\to\nu_{\mu}}$ $=\sin^{2}(2\theta)\,\sin^{2}\left(\frac{\Delta\Phi}{2}\right)$ (29)
@@
  • Both mass eigenstates can be spin-up:
@@
  • Both mass eigenstates can be spin-down:
@@
  • $\nu_{2}$ can be spin-up while $\nu_{1}$ can be spin-down:
@@
  • $\nu_{2}$ can be spin-down while $\nu_{1}$ can be spin-up:
@@
    Spin-dependent geometric phase: The parameters $a_{2}$ and $a_{4}$ contribute only when the mass eigenstates have opposite helicities. This asymmetric contribution, positive for more configuration and negative for the other, averages to zero for unpolarized neutrino beams but could manifest in polarization-sensitive experiments.
@@ Acknowledgments
This study was supported by Scientific and Technological Research Council of Türkiye (TUBITAK) under Grant Number 124F325. The authors thank TUBITAK for their support. The authors sincerely thank the anonymous referees for their thorough and attentive review of the manuscript, and for their insightful suggestions and constructive criticisms, which have significantly contributed to improving the clarity and overall quality of the paper.